# Exclusion of a direct progenitor detection for the Type Ic SN 2017ein based on late-time observations

### v1

@@ Abstract
To date, SN 2017ein is the only Type Ic supernova with a directly identified progenitor candidate. This candidate points to a very massive ($>$45 $M_{\odot}$) Wolf-Rayet progenitor, but its disappearance after the explosion of SN 2017ein remains unconfirmed. In this work, we revisit SN 2017ein in late-time images acquired by the Hubble Space Telescope (HST) at 2.4–3.8 yrs after peak brightness. We find this source has not disappeared and its brightness and color remain almost the same as in the pre-explosion images. Thus, we conclude that the pre-explosion source is not the genuine progenitor of SN 2017ein. We exclude the possibility that it is a companion star of the progenitor, since it has a much lower extinction than SN 2017ein; its color is also inconsistent with a star cluster, indicated by the newly added magnitude limit in F336W, apart from F [^5] W and F814W. We suggest, therefore, this source is an unrelated star in chance alignment with SN 2017ein. Based on the low ejecta mass, we propose that SN 2017ein is most likely originated from a moderately massive star with $M_{\rm ini}$ $\sim$ 8–20 $M_{\odot}$, stripped by binary interaction, rather than a very massive Wolf-Rayet progenitor.
@@ 1 Introduction
The difficulty in direct progenitor detections motivates researchers to use alternative methods to determine the dominant evolutionary channel for the Type Ib/Ic SN progenitors. For example, Smith et al. (2011a) argued that the occurrence rate of Type Ib/Ic SNe is too high to be compatible with the small number of massive WR stars at the high-mass end of the stellar initial mass function. Lyman et al. (2016), Taddia et al. (2018) and Woosley et al. (2021) analyzed the light curves of stripped-envelope SNe (SESNe; Types IIb, Ib and Ic), all supporting interacting binaries as the dominant progenitor channel for Type Ib/Ic SNe. By analyzing the nebular-phase spectra and SN environments, Fang et al. (2018) and Sun et al. (2023) suggested that Type Ic SN progenitors may undergo a hybrid envelope-stripping mechanism, in which the hydrogen envelopes are stripped by binary interaction while the helium envelopes are stripped by intense stellar winds.
@@ 2 Data and photometry
The data used in this work are from observations conducted with the HST (see Table 1 for a complete list). The pre-explosion images were observed with the Wide Field Planetary Camera 2 (WFPC2) Planetary Camera (PC), while the late-time images were acquired with the Wild Field Camera 3 (WFC3) Ultraviolet-Visible (UVIS) channel at t = 2.4–3.8 years after the peak brightness. All the observations were conducted in the F [^5] W and F814W filters, with an addition of F336W at t = 2.5 yrs. Besides, an F438W image, taken right after the SN explosion, was used as a reference image for differential astronomy.
We retrieved the data from the Barbara A. Mikulski Archive for Space Telescopes (MAST; https://mast.stsci.edu). For the pre-explosion WFPC2 images, we used the iraf task crrej (Gonzaga 2002) to remove cosmic rays. For the late-time WFC3 images, we used the tweakreg (Avila et al. 2014) package to align the individual exposures (*_flc.fits images) and combined them with the astrodizzle (Hack et al. 2012) package to remove the cosmic rays.
@@ 3 Results and discussion
Figure 1 shows the site of SN 2017ein in the pre-explosion and late-time observations. For better determination of the SN position, differential astrometry was used to align the images with respect to the image taken near peak brightness. We selected more than ten common stars as reference stars, and the accuracy of the alignment reaches 0.7 pixel. It is clearly shown that in the pre-explosion images, there is a source at the SN position in both F [^5] W and F814W bands, with a positional difference of only $\sim$0.13 pixel. This source corresponds to the progenitor candidate of SN 2017ein as reported in the previous studies (K18, VD18, and X19). For convenience, we shall refer to it as Source A hereafter. In the late-time images, however, Source A does not disappear and the SN position remains still bright.
@@
According to these phenomena, the most likely scenario is that the pre-explosion object, a star or star cluster, has not disappeared after the explosion of SN 2017ein since the late-time brightness and color are very consistent with the pre-explosion ones. The extra light in F814W and the red color at $t$ = 2.5 yrs could arise from a light echo (Boffi et al. 1999; Bond et al. 2003) or CSM interaction (Pellegrino et al. 2022; Fraser 2020), which has completely faded by the last epoch. A light echo or CSM interaction could also be present at $t$ = 2.4 yrs, but the observations have such large photometric uncertainties that we could not identify their contribution.
@@
We analyze Source A on the (F336W--F [^5] W, F [^5] W--F814W) color-color diagram (Figure 3) to determine whether it is a star or a star cluster. To reduce the photometric uncertainties in F [^5] W/F814W, we use the inverse-variance weighted averages of the pre-explosion magnitude and that at the last epoch as the adopted magnitude of Source A (i.e. 24.73 $\pm$ 0.06 mag and 24.69 $\pm$ 0.11 mag, respectively). For F336W, we conservatively regard the derived magnitude of 23.61 $\pm$ 0.06 mag as a brightness upper limit for the underlying star or star cluster, since this magnitude is measured at t = 2.5 yrs, by which time there could still be some extra light caused by a light echo or CSM interaction. For comparison, we show synthetic colors for stars of different effective temperatures (based on the Castelli & Kurucz 2004 stellar atmospheric models) and star clusters of different ages (based on the bpass model population synthesis; Eldridge et al. 2017; Stanway & Eldridge 2018; Byrne et al. 2022). For both stellar and cluster SEDs, we also apply a range of different extinctions.
In Fig. 3(a), the color of Source A appears very different from that expected for star clusters. The previous studies (K18, VD18, and X19) also discussed the possibility of Source A being a young star cluster but only with the F [^5] W -- F814W color from the pre-explosion observations. Now with the addition of the F336W limit, we suggest this scenario is not very likely since the young star clusters would be too bright in this filter. This analysis, however, is based on a standard extinction law of $R_{V}$ = 3.1; the color of Source A could still be consistent with that of a star cluster in case of a different host-galaxy extinction law.
In Fig. 3(b), the color of Source A appears consistent with those of stellar models with effective temperature of $T_{\rm eff}\lesssim$ 16000–21000 K and (total) extinction of $A_{V}<0.23$ mag, which suggests that Source A is most likely to be a hot massive star. However, it is not the progenitor star of SN 2017ein since it has not vanished after explosion. It is worth noting that the derived total extinction of Source A is much smaller than that of SN 2017ein itself (1.26 $\pm$ 0.2 mag, K18; 1.0–1.9 mag, VD18; 1.29 $\pm$ 0.18 mag, X19; measured with different methods). We argue therefore that it is not a companion star of SN 2017ein but an unrelated star in chance alignment.
@@ 4 Summary
In this work, we re-investigate the progenitor candidate reported for SN 2017ein with the new HST images acquired at 2.4–3.8 yrs after peak brightness. We find that the candidate has not disappeared, with its brightness and color almost the same as the pre-explosion ones. Thus, we conclude that the candidate is not the genuine progenitor. It is neither a companion star of the progenitor considering its much lower extinction than SN 2017ein. The color-color diagram analysis also rules out the possibility of this source being a star cluster. Therefore, we suggest it is most likely to be an unrelated star in chance alignment with SN 2017ein. Based on the low ejecta mass, we suggest that the progenitor is more likely to be a moderately massive star stripped in a binary, rather than a very massive WR star as suggested by previous works.
@@ acknowledgments
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB0550300. NCS’s research is funded by the NSFC grants No. 12303051 and No. 12261141690 and ZXN acknowledges support from the NSFC through grant No. 12303039. JFL acknowledges support from the NSFC through grants No. 11988101 and No. 11933004 and from the New Cornerstone Science Foundation through the New Cornerstone Investigator Program and the XPLORER PRIZE.

### v2

@@ Abstract
To date, SN 2017ein is the only Type Ic supernova with a directly identified progenitor candidate. This candidate points to a very massive ($>$45 $M_{\odot}$) Wolf-Rayet progenitor, but its disappearance after the explosion of SN 2017ein remains unconfirmed. In this work, we revisit SN 2017ein in late-time images acquired by the Hubble Space Telescope (HST) at 2.4–3.8 yrs after peak brightness. We find this source has not disappeared and its brightness and color remain almost the same as in the pre-explosion images. Thus, we conclude that the pre-explosion source is not the genuine progenitor of SN 2017ein. It is not much likely to be a companion star of the progenitor, since it has a much lower extinction than SN 2017ein; its color also seems inconsistent with a star cluster, indicated by the newly added magnitude limit in F336W, apart from F [^5] W and F814W. We suggest, therefore, this source is an unrelated star in chance alignment with SN 2017ein. Based on the low ejecta mass, we propose that SN 2017ein is most likely originated from a moderately massive star with $M_{\rm ini}$ $\sim$ 8–20 $M_{\odot}$, stripped by binary interaction, rather than a very massive Wolf-Rayet progenitor.
@@ 1 Introduction
The difficulty in direct progenitor detections motivates researchers to use alternative methods to determine the dominant evolutionary channel for the Type Ib/Ic SN progenitors. For example, Smith et al. (2011a) argued that the occurrence rate of Type Ib/Ic SNe is too high to be compatible with the small number of massive WR stars at the high-mass end of the stellar initial mass function. Lyman et al. (2016), Taddia et al. (2018) and Woosley et al. (2021) analyzed the light curves of stripped-envelope SNe (SESNe; Types IIb, Ib and Ic), all supporting interacting binaries as the dominant progenitor channel for Type Ib/Ic SNe. By analyzing the nebular-phase spectra and SN environments, Fang et al. (2018) and Sun et al. (2023) suggested that Type Ic SN progenitors may undergo a hybrid envelope-stripping mechanism, in which the hydrogen envelopes are stripped by binary interaction while the helium envelopes are stripped by intense stellar winds. Solar et al. (2024) confirmed observationally for a statistically large sample that Type Ic SN progenitors are binary systems, using molecular clouds in the environments of the SN explosions as tracer of lifetimes.
@@ 2 Data and photometry
The data used in this work are from observations conducted with the HST (see Table 1 for a complete list). The pre-explosion images were observed with the Wide Field Planetary Camera 2 (WFPC2) Planetary Camera (PC), while the late-time images were acquired with the Wild Field Camera 3 (WFC3) Ultraviolet-Visible (UVIS) channel at t = 2.4–3.8 years after the peak brightness. All the observations were conducted in the F [^5] W and F814W filters, with an addition of F336W at t = 2.5 yrs. Besides, an F438W image, taken right after the SN explosion, was used as a reference image for differential astronomy. The HST data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/7n8j-cy13 (catalog https://doi.org/10.17909/7n8j-cy13).
We retrieved the data from the Barbara A. Mikulski Archive for Space Telescopes (MAST; https://mast.stsci.edu). For the pre-explosion WFPC2 images, we used the iraf task crrej (Gonzaga 2002) to remove cosmic rays. For the late-time WFC3 images, we used the tweakreg (Avila et al. 2015) package to align the individual exposures (*_flc.fits images) and combined them with the astrodizzle (Hack et al. 2012) package to remove the cosmic rays.
@@ 3 Results and discussion
Figure 1 shows the site of SN 2017ein in the pre-explosion and late-time observations. For better determination of the SN position, differential astrometry was used to align the images with respect to the image taken near peak brightness. We selected more than ten common stars as reference stars, and the uncertainty of the alignment reaches 0.7 pixel. It is clearly shown that in the pre-explosion images, there is a source at the SN position in both F [^5] W and F814W bands, with a positional difference of only $\sim$0.13 pixel. This source corresponds to the progenitor candidate of SN 2017ein as reported in the previous studies (K18, VD18, and X19). For convenience, we shall refer to it as Source A hereafter. In the late-time images, however, Source A does not disappear and the SN position remains still bright. The sharpness parameter of Source A reported by dolphot is very close to zero, which implies that Source A is a point source.
@@
According to these phenomena, the most likely scenario is that the pre-explosion object, a star or star cluster, has not disappeared after the explosion of SN 2017ein since the late-time brightness and color are very consistent with the pre-explosion ones. The extra light in F814W and the red color at $t$ = 2.5 yrs could arise from a light echo (Boffi et al. 1999; Bond et al. 2003) or CSM interaction (Pellegrino et al. 2022; Fraser 2020), which has completely faded by the last epoch. A light echo or CSM interaction could also be present at $t$ = 2.4 yrs, but the observations have such large photometric uncertainties that we could not identify their contribution. It is noteworthy that another scenario is also considered, where the SN ejecta shocks the surviving binary companion, causing it to be over luminous and inflated for several years after the SN explosion (Hirai et al., 2018; Ogata et al., 2021; Chen et al., 2023). However, in such a scenario, there is little possibility that Source A’s magnitudes of different bands remain this consistent with the pre-explosion ones.
@@
We analyze Source A on the (F336W--F [^5] W, F [^5] W--F814W) color-color diagram (Figure 3) to determine whether it is a star or a star cluster. To reduce the photometric uncertainties in F [^5] W/F814W, we use the inverse-variance weighted averages of the pre-explosion magnitude and that at the last epoch as the adopted magnitude of Source A (i.e. 24.73 $\pm$ 0.06 mag and 24.69 $\pm$ 0.11 mag, respectively). For F336W, we conservatively regard the derived magnitude of 23.61 $\pm$ 0.06 mag as a brightness upper limit for the underlying star or star cluster, since this magnitude is measured at t = 2.5 yrs, by which time there could still be some extra light caused by a light echo or CSM interaction. This possible extra light may also influence the brightness in F [^5] W and F814W, but the effects are believed to be negligible (see above discussion). For comparison, we show synthetic colors for stars of different effective temperatures (based on the Castelli & Kurucz 2004 stellar atmospheric models) and star clusters of different ages (based on the bpass model population synthesis; Eldridge et al. 2017; Stanway & Eldridge 2018; Byrne et al. 2022). For both stellar and cluster SEDs, we also apply a range of different extinctions.
In Fig. 3(a), the color of Source A appears very different from that expected for star clusters. The previous studies (K18, VD18, and X19) also discussed the possibility of Source A being a young star cluster but only with the F [^5] W -- F814W color from the pre-explosion observations. Now with the addition of the F336W limit, we suggest this scenario is not very likely since the young star clusters would be too bright in this filter. This analysis, however, is based on a standard extinction law of $R_{V}$ = 3.1; the color of Source A could still be consistent with that of a star cluster in case of a different host-galaxy extinction law. Moreover, since Source A lies quite close to the cluster models on the diagram, we might also consider the possibility of it truly being a cluster, with an age of $10^{6.5}$ yrs and a total extinction of $A_{V}=0.3$ mag. We estimated the mass of such a cluster to be $10^{3}$ $M_{\odot}$. Binaries or single stars with $M<114M_{\odot}$ can explode at this age, based on BPASS models. It is noteworthy that this extinction value is much lower than that of SN 2017ein (1.26 $\pm$ 0.2 mag, K18; 1.0–1.9 mag, VD18; 1.29 $\pm$ 0.18 mag, X19; measured with different methods), which might be explained by the supposition that the SN explodes from a dusty star within the cluster. Moreover, we find the influence of metallicity on the cluster models is negligible in our work.
[…]
@@ 4 Summary
In this work, we re-investigate the progenitor candidate reported for SN 2017ein with the new HST images acquired at 2.4–3.8 yrs after peak brightness. We find that the candidate has not disappeared, with its brightness and color almost the same as the pre-explosion ones. Thus, we conclude that the candidate is not the genuine progenitor. It is neither a companion star of the progenitor considering its much lower extinction than SN 2017ein. The color-color diagram analysis also indicates that this source is less likely to be a star cluster. Therefore, we suggest it is most likely to be an unrelated star in chance alignment with SN 2017ein. Based on the low ejecta mass, we suggest that the progenitor is more likely to be a moderately massive star stripped in a binary, rather than a very massive WR star as suggested by previous works.
@@ acknowledgments
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB0550300. NCS’s research is funded by the NSFC grants No. 12303051 and No. 12261141690 and ZXN acknowledges support from the NSFC through grant No. 12303039. JFL acknowledges support from the NSFC through grants No. 11988101 and No. 11933004 and from the New Cornerstone Science Foundation through the New Cornerstone Investigator Program and the XPLORER PRIZE. We thank the anonymous referee for the helpful comments.
@@ Appendix A Systematic error correction
Figure A2:  Histograms of magnitude differences with respect to the epoch of t = 3.8 yrs. The solid lines are Gaussian fit curves and the dashed lines represent the means of the Gaussian fit. Normalization has been applied to each graph.